OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (5, 0) and a directrix at x = -5.

OpenStudy (anonymous):

@helpme1.2 help??

OpenStudy (anonymous):

standard form of parabola = ax^2 + bx + c

OpenStudy (anonymous):

so a25-5x+0?

OpenStudy (anonymous):

@helpme1.2

OpenStudy (anonymous):

x+5 = sqrt((x-5)^2 + (y-0)^2)

OpenStudy (anonymous):